OpenStudy (anonymous):

what is the simplified form of the (sqrt) k17

OpenStudy (king):

\[\sqrt{17k}?\]

OpenStudy (anonymous):

\[\sqrt{k^17}=\sqrt{k*k^{16}}=k^8\sqrt{k}\]

OpenStudy (anonymous):

thank you

OpenStudy (king):

is it \[\sqrt{k ^{17}}\]?

OpenStudy (anonymous):

yes